### v2

@@ 3 Numerical results
The truncation is performed in order to have an initial condition with compact support; the form of the equations is such that the solution will still be compactly supported at all times [aronson80, newman80]. In all the simulations we set $\boldsymbol{x}_{0}=(0,0)$, $y_{0}=\varphi(O(0,\boldsymbol{x}_{0}))$, $\theta_{\boldsymbol{x}}=0.5$, $\theta_{y}=0.5$. The parameter $A_{u}$ is set to
@@
In the particular situation of a circle, this definition coincides with the standard diameter; in the general case, the diameter is the longest length found inside the set. This choice is based on the assumption that small tumours cannot be clinically detected, hence the therapy may only start when cancer cells reach a density of at least one tenth of the carrying capacity in a large region. We set $d_{\text{inj}}=5.2\;$mm, as in Ref. [kim06] and the central viral injection takes the form
@@
with $A_{v}=7\times 10^{9}$, $\theta_{v}=0.5$. This allows for a total number of viral particles in agreement with the experiments performed in Ref. [kim06].
@@ 3.3.2 Hypoxia-specific oncolytic viruses (case 3b)
All the simulations presented so far rely on the assumption that viral infection is less effective in hypoxic cells: this can be explained by their slower metabolic activity, which also affects the translation of viral proteins [shengguo11]. Nevertheless, one should note that some particular oncolytic viruses can specifically target receptors that are upregulated in case of the lack of oxygen [sadri23, shengguo11]. This property appears particularly promising in light of the ineffectiveness of most classic cancer therapies in hypoxic conditions [zhuang23].
@@
Although the use of hypoxia-selective oncolytic viruses appears promising, our findings also stress that tumour eradication is still far from achieved. Furthermore, the cell lines selected are the most proliferative; hence, if the viral infection stopped for some external reason (such as immune response), the tumour would quickly regain its aggressiveness. Nonetheless, it is essential to observe that the average epigenetic trait is significantly reduced with respect to the one that environmental conditions would select; consequently, the tumour is now sensitive to standard treatments that lose their effectiveness in hypoxic conditions. This suggests that a hypoxic tumour could be effectively treated using a combination of therapies targeting cells with different degrees of adaptation to hypoxia [shengguo11].
@@ 4 Conclusions
The key assumption of this work is that, in general, the slower metabolic activity of hypoxic cells is associated with a less effective viral infection. This has been modelled by considering an infection rate $\beta$ that depends on the adaptation of cells to hypoxia. Our choice is motivated by the existing literature [lorenzi21] and the possibility of obtaining some theoretical insights. Other approaches would also be possible: for example, one may directly consider the dependence of viruses on the host translational machinery to translate viral proteins; this would correspond to a viral burst size $\alpha$ that depends on the oxygen concentration or on the level of metabolic activity of the cell (which could correspond to the variable $y$).
There are several promising directions for future research and the results may be extended in several ways, both from the mathematical and the modeling perspectives. From the mathematical point of view, the formal asymptotic analysis of A may constitute a good starting point for characterising the travelling waves shown in the numerical simulations, at least in simplified settings. From a biological point of view, we aim to develop a model to study the interaction of oncolytic virotherapy and radiotherapy, which aligns with the clinical interest of combining the two therapies. Radiotherapy is well-known to decrease its effectiveness in hypoxic conditions and several mathematical models similar to the one in the present work have been developed to investigate this phenomenon [celora21, chiari23radio]. Although the combination with oncolytic viruses that decrease their efficacy due to hypoxia does not appear beneficial, using hypoxia-targeting oncolytic viruses could be promising.
@@ Appendix A Formal asymptotic analysis
It is important to note that the system may not converge to an equilibrium: indeed, for some parameter values, central oscillations persist even for very long times, similar to the ones described in Refs. [baabdulla23, pooladvand21] (see also electronic supplementary material S9).
@@
Under the concavity hypotheses explained in Section 2.2, we expect $n_{\varepsilon}$ to be a strictly concave function of $y$; this is also true for $n$ [mirrahimi2015asymptotic, perthame08]. We then define
@@
As a consequence, unlimited growth of cancer cells is not possible, i.e. $\rho_{\varepsilon}(t)<+\infty$ for all $\varepsilon$ (see also, for example, Ref. [lorenzi21invasion]). Therefore,
@@ B.1 Parameter values
The spatial diffusion coefficient of tumour cells $D_{\boldsymbol{x}}$ has been estimated from the experimental data of the U343 control group of [kim06], as already done in Refs. [morselli23, pooladvand21]. In these experiments, the tumour volume passes in 40 days from 7070\;mm3 to 10001000\;mm3, which corresponds to a change in the tumour radius from approximately 2.62.6\;$m$ to approximately 6.26.2\;$m$. Hypoxia may play some role in the process, but this is not taken into account in their data: for the sake of simplicity, we assume a moderately hypoxic situation so that the proliferation rate takes the value $p(0.5)$ (i.e., the average between $p_{M}$ and $p_{m}$). Hence, the dynamics of uninfected cells in the absence of viral infection follow the equation
@@
As we mention in the main text, an initial condition with compact support evolves into a wave that travels with the minimal speed $\sqrt{D_{\boldsymbol{x}}Kp(0.5)/2}$ [aronson80, newman80]; this yields the estimate
@@
The oxygen diffusion coefficient $D_{O}$ has been set to 3.603.60\;mm2/h, as in Ref. [mueller-klieser84].
@@ B.2 Numerical method
Numerical simulations use a finite volume method developed by adapting the procedures presented in Refs. [bessemoulin12, carrillo15] to our problem. We discretise the space $\Omega$ with a uniform mesh consisting of the cells

### v3

@@ 3 Numerical results
The truncation is performed in order to have an initial condition with compact support; the form of the equations is such that the solution will still be compactly supported at all times [7, 67]. In all the simulations we set $\boldsymbol{x}_{0}=(0,0)$, $y_{0}=\varphi(O(0,\boldsymbol{x}_{0}))$, $\theta_{\boldsymbol{x}}=0.5$, $\theta_{y}=0.5$. The parameter $A_{u}$ is set to
@@
In the particular situation of a circle, this definition coincides with the standard diameter; in the general case, the diameter is the longest length found inside the set. This choice is based on the assumption that small tumours cannot be clinically detected, hence the therapy may only start when cancer cells reach a density of at least one tenth of the carrying capacity in a large region. We set $d_{\text{inj}}=5.2\;$mm, as in Ref. [49] and the central viral injection takes the form
@@
with $A_{v}=7\times 10^{9}$, $\theta_{v}=0.5$. This allows for a total number of viral particles in agreement with the experiments performed in Ref. [49].
@@ 3.3.2 Hypoxia-specific oncolytic viruses (case 3b)
All the simulations presented so far rely on the assumption that viral infection is less effective in hypoxic cells: this can be explained by their slower metabolic activity, which also affects the translation of viral proteins [81]. Nevertheless, one should note that some particular oncolytic viruses can specifically target receptors that are upregulated in case of the lack of oxygen [79, 81]. This property appears particularly promising in light of the ineffectiveness of most classic cancer therapies in hypoxic conditions [94].
@@
Although the use of hypoxia-selective oncolytic viruses appears promising, our findings also stress that tumour eradication is still far from achieved. Furthermore, the cell lines selected are the most proliferative; hence, if the viral infection stopped for some external reason (such as immune response), the tumour would quickly regain its aggressiveness. Nonetheless, it is essential to observe that the average epigenetic trait is significantly reduced with respect to the one that environmental conditions would select; consequently, the tumour is now sensitive to standard treatments that lose their effectiveness in hypoxic conditions. This suggests that a hypoxic tumour could be effectively treated using a combination of therapies targeting cells with different degrees of adaptation to hypoxia [81].
@@ 4 Conclusions
The key assumption of this work is that, in general, the slower metabolic activity of hypoxic cells is associated with a less effective viral infection. This has been modelled by considering an infection rate $\beta$ that depends on the adaptation of cells to hypoxia. Our choice is motivated by the existing literature [59] and the possibility of obtaining some theoretical insights. Other approaches would also be possible: for example, one may directly consider the dependence of viruses on the host translational machinery to translate viral proteins; this would correspond to a viral burst size $\alpha$ that depends on the oxygen concentration or on the level of metabolic activity of the cell (which could correspond to the variable $y$).
There are several promising directions for future research and the results may be extended in several ways, both from the mathematical and the modeling perspectives. From the mathematical point of view, the formal asymptotic analysis of A may constitute a good starting point for characterising the travelling waves shown in the numerical simulations, at least in simplified settings. From a biological point of view, we aim to develop a model to study the interaction of oncolytic virotherapy and radiotherapy, which aligns with the clinical interest of combining the two therapies. Radiotherapy is well-known to decrease its effectiveness in hypoxic conditions and several mathematical models similar to the one in the present work have been developed to investigate this phenomenon [19, 21]. Although the combination with oncolytic viruses that decrease their efficacy due to hypoxia does not appear beneficial, using hypoxia-targeting oncolytic viruses could be promising.
@@ Appendix A Formal asymptotic analysis
It is important to note that the system may not converge to an equilibrium: indeed, for some parameter values, central oscillations persist even for very long times, similar to the ones described in Refs. [9, 74] (see also electronic supplementary material S9).
@@
Under the concavity hypotheses explained in Section 2.2, we expect $n_{\varepsilon}$ to be a strictly concave function of $y$; this is also true for $n$ [63, 70]. We then define
@@
As a consequence, unlimited growth of cancer cells is not possible, i.e. $\rho_{\varepsilon}(t)<+\infty$ for all $\varepsilon$ (see also, for example, Ref. [54]). Therefore,
@@ B.1 Parameter values
The spatial diffusion coefficient of tumour cells $D_{\boldsymbol{x}}$ has been estimated from the experimental data of the U343 control group of [49], as already done in Refs. [65, 74]. In these experiments, the tumour volume passes in 40 days from 7070\;mm3 to 10001000\;mm3, which corresponds to a change in the tumour radius from approximately 2.62.6\;$m$ to approximately 6.26.2\;$m$. Hypoxia may play some role in the process, but this is not taken into account in their data: for the sake of simplicity, we assume a moderately hypoxic situation so that the proliferation rate takes the value $p(0.5)$ (i.e., the average between $p_{M}$ and $p_{m}$). Hence, the dynamics of uninfected cells in the absence of viral infection follow the equation
@@
As we mention in the main text, an initial condition with compact support evolves into a wave that travels with the minimal speed $\sqrt{D_{\boldsymbol{x}}Kp(0.5)/2}$ [7, 67]; this yields the estimate
@@
The oxygen diffusion coefficient $D_{O}$ has been set to 3.603.60\;mm2/h, as in Ref. [66].
@@ B.2 Numerical method
Numerical simulations use a finite volume method developed by adapting the procedures presented in Refs. [13, 18] to our problem. We discretise the space $\Omega$ with a uniform mesh consisting of the cells